OpenStudy (anonymous):

p + 1/3p = 2/3 solve for p.

OpenStudy (astrophysics):

p+1/3p is the same thing 1+1/3 just with units, so do that first.

OpenStudy (anonymous):

the answer is 1/2 on my textbook but i need to answer it step by step

OpenStudy (astrophysics):

What's 1+1/3?

OpenStudy (anonymous):

1 1/2?

OpenStudy (anonymous):

3p+1/P=2/3 9p+3=2p 9p-2p=-3 7p=-3 p=-3/7 ans.

OpenStudy (astrophysics):

That's wrong... @santanu74

OpenStudy (astrophysics):

@xLogic What's 1+1/3 = ?

OpenStudy (astrophysics):

Think of 1 as a fraction so \[\frac{ 1 }{ 1 }+\frac{ 1 }{ 3 } =?\]

OpenStudy (anonymous):

2/3?

OpenStudy (astrophysics):

I don't know how you got that, but it seems you're weak on fractions, before you go to this problem, please review how to do fractions.

OpenStudy (astrophysics):